# Effects of Turbulent Energy Exchange between Electrons and Ions on Global Temperature Profiles

### v1

@@ Abstract
In the present work, we extend those previous investigations of the local effects of turbulent energy exchange and examine its impact on the steady-state global temperature profiles. To predict the global temperature profiles, we employ GOTRESS[5, 6], which is a one-dimensional steady-state transport solver to solve the heat transport equations at a steady state given by,
@@
In this study, the impact of turbulent energy exchange on steady-state temperature profiles is investigated. As long as the heating powers for ions and electrons do not differ drastically (e.g., by less than a factor of two), the contribution of turbulent energy exchange remains small. However, when large disparities in the heating or energy fluxes of electrons and ions occur, turbulent energy exchange becomes comparable to, or even exceeds, the collisional contribution, substantially altering the temperature profiles. These results highlight scenarios in which turbulent energy exchange must be incorporated for accurate reactor-scale predictions. It is worthwhile to investigate the impact of turbulent energy exchange in plasma start-up scenarios, where electron heating is dominated by Ohmic heating and ions are heated solely through energy transfer from electrons.[19].
The present study is supported in part by the JSPS Grants-in-Aid for Scientific Research Grant No. 24K07000 and in part by the NIFS Collaborative Research Program NIFS23KIPT009. This work is also supported by JST SPRING, Grant Number JPMJSP2108. Simulations in this work were performed on “Plasma Simulator” (NEC SX-Aurora TSUBASA) of NIFS with the support and under the auspices of the NIFS Collaboration Research program (NIFS24KISM007).

### v2

@@ Abstract
In the present work, by comparing simulation results with and without the turbulent energy exchange term, we examine its impact on the steady-state global temperature profiles and investigate the necessity of incorporating turbulent energy exchange into the analysis of existing experiments and the prediction of global temperature profiles for future fusion reactors. To predict the global temperature profiles, we employ GOTRESS[5, 6], which is a one-dimensional steady-state transport solver to solve the heat transport equations at a steady state given by,
@@
In this study, the impact of turbulent energy exchange on steady-state temperature profiles is investigated. As long as the heating powers for ions and electrons do not differ drastically (e.g., by less than a factor of two), the contribution of turbulent energy exchange remains small. However, when large disparities in the heating or energy fluxes of electrons and ions occur, turbulent energy exchange becomes comparable to, or even exceeds, the collisional contribution, substantially altering the temperature profiles. These results highlight scenarios in which turbulent energy exchange must be incorporated for accurate reactor-scale predictions. It is worthwhile to investigate the impact of turbulent energy exchange in plasma start-up scenarios, where electron heating is dominated by Ohmic heating and ions are heated solely through energy transfer from electrons.[19]. As a future task, it remains to validate the results using the MMM-calibrated BgB model by performing gyrokinetic simulations over the entire minor radius. In addition, since the plasma start-up phases are inherently non-steady-state, further investigation is required to properly address this regime.
The present study is supported in part by the JSPS Grants-in-Aid for Scientific Research Grant No. 24K07000 and in part by the NIFS Collaborative Research Program NIFS23KIPT009. This work is also supported by JST SPRING, Grant Number JPMJSP2108. Simulations in this work were performed on “Plasma Simulator” with the support and under the auspices of the NIFS Collaboration Research program (NIFS24KISM007).